# Promise and Peril: Stellar Contamination and Strict Limits on the Atmosphere Composition of TRAPPIST-1c from JWST NIRISS Transmission Spectra

### v1

@@ 3 Modelling of the transit light source effect
We first model the transmission spectrum of TRAPPIST-1 c assuming that the TLSE can completely explain each visit’s transmission spectrum. This approach is motivated by the presence of small features resembling water absorption (e.g., around 1.4 µm; Figure 2) and a pronounced slope towards short wavelengths in the spectrum of the first visit. We use the open-source code stctm (Piaulet-Ghorayeb, 2024) to perform the TLS retrievals. We model the presence of two populations of heterogeneities: spots (covering a fraction of the photosphere, and cooler than the photosphere at a temperature ), and faculae (with representing their covering fraction, and their temperature).
@@
For each population of heterogeneities, we fit for the temperature difference to the photosphere (), and place a conservative uniform prior from 0 to 50% on the covering fraction for each of the components, . The stellar photosphere temperature is fitted using a Gaussian prior, with the mean set by the literature value of 2566 K (Agol et al., 2021) and a standard deviation of 50 K. We allow faculae to be at most 1000 K hotter than the photosphere (with an otherwise uniform prior on ), while the minimum temperature of spots is set by the minimum temperature for which stellar models are available in our grid (2300 K). In the retrieval, we select at each iteration the closest-matching model (in effective temperature and surface gravity space) for the photosphere, spot, and faculae components from a finely-sampled pre-computed grid obtained with the MSG module (Townsend & Lopez, 2023) using the PHOENIX (Husser et al., 2013) stellar model grid. We also fit the used for the heterogeneities, allowing it to be lower than that of the photosphere following recent work (Lim et al., 2023; Fournier-Tondreau et al., 2024), but retrieve a value consistent with that of the photosphere (Agol et al., 2021).
@@
We use the affine-invariant Markov Chain Monte Carlo sampler emcee (Foreman-Mackey et al., 2013) to explore the parameter space, with the number of walkers set as 20 times the number of fitted parameters, and run the chains for 5000 steps (60% are discarded as burn-in). In addition to the stellar heterogeneity properties, we fit the wavelength-independent transit depth of TRAPPIST-1 c, . We perform independent fits to the spectrum of each visit, as they are separated by about one year and should not, a priori, be affected by the same TLS contribution. We find that similar spot and facula properties can explain both visits (Figure 2).
@@ 4 Joint modelling of atmosphere and stellar contamination
We perform three categories of retrievals corresponding to different assumptions about the planetary atmosphere: (1) H2/He-dominated scenarios; (2) single-component high-mean molecular weight atmospheres; (3) two-component high-mean molecular weight atmospheres with one infrared absorber present in the atmosphere and either N2 or O2 acting as the background gas.
@@
We parameterize H2/He-dominated atmospheres in the retrieval with a H2/He mixture as the background gas (assuming He/H2= 0.157), log-uniform priors on the abundances of N2, CH4, H2O, CO, CO2, NH3, and SO2, and a log-uniform prior on the pressure at which the atmosphere becomes opaque (either due to the presence of a cloud deck or a solid surface).
@@
For single-component high-mean molecular weight atmospheres, we only explore absorbers with prominent features over the NIRISS/SOSS wavelength range; namely, atmospheres made of 100% CH4, H2O, CO, or NH3. We also fit for the “effective surface pressure” at which the atmosphere becomes opaque.
@@
Finally, we also explore potential N2-dominated atmospheres (similar to some terrestrial objects in the solar system), or O2-dominated atmospheres (motivated by the possible O2-CO2 atmosphere consistent with the MIRI eclipse observations of TRAPPIST-1 c; (Zieba et al., 2023)) where only one of CH4, H2O, CO, or NH3 acts as an absorber, beyond the broad N2-N2 or O2-O2 collision-induced absorption features introduced by the background gas. In this retrieval setup, we fit for the partial pressure of each component.
@@
We assume that the vertical temperature structure is isothermal, motivated by the likely-thin region of the upper atmosphere probed by our observations in transmission, and only fit for this photosphere temperature. We use the nested sampling method (Skilling, 2004, 2006) to sample the full parameter space, with the multi-ellipsoid method implemented using the nestle module 222https://github.com/kbarbary/nestle within the SCARLET retrieval framework. The final posterior distributions are obtained from retrievals performed using at least 1000 live points. We compute forward models at R=15,625 before convolving them to the resolution of the data assuming a uniform throughput within each bin for the likelihood calculation.
@@
We also calculate forward models for 1 and 100 solar metallicity, H2/He-dominated atmospheres with varying cloud-top pressures (Figure 3; left panel), and models for several high mean molecular weight compositions (Figure 4; top panel), to put our retrieval results in context. For each model, the temperature-pressure profile is assumed to be isothermal at the equilibrium temperature of TRAPPIST-1 c.
For each set of retrievals, we account in the modelling for the fact that the transmission spectra we observe are likely heavily affected by stellar contamination (see Figure 2), thereby marginalizing over the interplay that unocculted stellar heterogeneities can have with planetary absorption features. In addition to the planetary atmosphere properties, we thus account for stellar heterogeneities assuming a single population of spots, fitting for the spot covering fraction and their temperature contrast with the photosphere, as well as the photosphere temperature itself. We also assume that the stellar heterogeneity components of both visits are identical, motivated by the results from the stctm fit (Figure 2), and do not fit for a different for the heterogeneity component. We note, though, that we performed one set of atmosphere+stellar contamination retrievals where we fit for visit-specific stellar heterogeneity properties, and found that this does not affect our inferences, further motivating our choice of shared spot parameters for both visits. We place uniform priors on the spot fraction and temperature, but use the same Gaussian prior as in the TLS-only fit for the photosphere temperature.
We note that we find a vertical offset of 100 ppm between the spectra of both visits (as can be inferred from the white light transit radii presented in Table 1). We attribute it to the presence of the flare in the pre-transit baseline of the second visit that may result in a dilution effect when fitting the transit light curves and therefore fit for a parameter describing the visit 1-visit 2 offset.
@@ 5 Discussion
5.1 Impacts of Stellar Contamination
Our two-transit study of TRAPPIST-1 c aptly summarizes the promise and peril of transmission observations of rocky planets around M dwarf host stars. There are now a plethora of studies that have shown that stellar contamination is a true bottleneck in the attempt to detect atmospheres around rocky, M dwarf planets via transmission spectroscopy (e.g., May et al., 2023; Moran et al., 2023; Lim et al., 2023; Cadieux et al., 2024). In most cases, different manifestations of the TLSE (i.e., different temperature distributions, covering fractions, etc. of stellar spots and faculae) give rise to obviously different transmission spectra between visits, allowing for a clear inference of the impacts of the TLSE, even without undertaking in-depth modelling. Consistency in the location and shapes of spectral features between visits has been argued to lend credence to an interpretation favouring a planetary atmosphere origin (e.g. May et al., 2023).
However, despite the high level of consistency between the transmission spectra (1.1- on average), our retrieval analysis still finds that both visits can be entirely explained by stellar contamination. As shown in Figure 2, we find very similar stellar heterogeneity parameters (spot/faculae temperatures and covering fractions) for both visits. The spot and faculae temperatures are consistent with those found by Lim et al. (2023) in their analysis of NIRISS/SOSS transmission spectra of TRAPPIST-1 b. However, we find both visits to be spot-dominated, with minimal contributions from faculae, whereas Lim et al. (2023) have one spot-dominated, and one facula-dominated visit.
The consistency between visits would perhaps be unsurprising if they were separated by a very small period of time, which would therefore probe similar spot and facula distributions on the host star. However, our two visits are separated by 367 d, significantly longer than the 3.3 d stellar rotation period (Luger et al., 2017). When phase-folded to the stellar rotation period, this corresponds to a phase difference of 0.21 — similar to the observations of Lim et al. (2023), who, as mentioned above, found no such consistency between their two visits. This highlights the fact that TLSE models are generally non-predictive of future stellar contamination, since the distributions of spots and faculae in the host star’s photosphere is entirely stochastic (Rackham et al., 2018, 2023; TRAPPIST-1 JWST Community Initiative et al., 2024).
@@ 5.2 Cloud-Free and Thick Hydrogen-Rich Atmospheres Are Ruled Out
We find that, similarly to the closer-in planet TRAPPIST-1 b (Lim et al., 2023), we can confidently rule out thick or cloud-free hydrogen-dominated atmospheres, with a 3- upper limit of only 0.1 bar on the effective surface pressure of a H2/He-dominated atmosphere atop TRAPPIST-1 c (Figure 3). The thin 0.1 bar hydrogen-rich atmosphere scenarios allowed by the transmission spectrum are, moreover, unlikely from an atmospheric evolution standpoint — TRAPPIST-1 c was exposed to large cumulative amounts of high-energy XUV irradiation from its host star, sufficient to strip away any primordial atmosphere (Hori & Ogihara, 2020; Turbet et al., 2020) even when considering hydrogen outgassing (Hu et al., 2023). The transmission spectrum still allows within the 3- confidence contours for thick, high-metallicity (100 solar) atmospheres, which may be more resilient against escape due to their higher mean molecular weights.
5.3 Limits on High Mean Molecular Weight Atmospheres
@@
We, therefore, turn to evaluate the range of plausible high-metallicity atmosphere scenarios for TRAPPIST-1 c, and focus on the detectability of H2O-, NH3-, CO- and CH4-rich atmospheres, as these infrared absorbers have sufficient opacity within the NIRISS/SOSS wavelength range to obtain meaningful constraints. We can rule out thick (1 bar) 100% H2O and NH3 atmospheres at better than 3, and even a 1 bar 100% CO atmosphere is excluded at the 2- level, although thinner 0.1 bar atmospheres remain plausible explanations for the flat spectrum (Figure 4).
@@
With the exception of CH4, we find tight upper limits on the abundance of all the other molecules tested (H2O, NH3, CO), even in lower-mean molecular weight atmospheres dominated by N2 or O2 as the background gas (Figure 5). We compare those limits, where available, with measured compositions of solar system planets. In particular, we highlight in Figure 5 the amount of H2O in Earth’s N2-dominated atmosphere (using 0.4% as the mean mole fraction over the entire 1-bar atmosphere, although it can reach about one percent at sea level; Kasai et al., 2011; Kelsey et al., 2022), or the amount of CH4 in Titan’s atmosphere, assuming either a cloud-free Titan scenario (where one can see down to the surface) with 5% CH4, or a hazy scenario accounting for the fact that hydrocarbon haze opacity raises the effective surface pressure to 10 mbar, and with a factor of three less methane in the upper atmosphere (see e.g. Charnay et al., 2014. We find that similar amounts of water to what is observed in Earth’s atmosphere are ruled out for TRAPPIST-1 c, while the transmission spectrum alone is not sensitive enough to the CH4 abundance to offer meaningful constraints on Titan-like atmospheres. Although our observations cannot rule out the presence of CH4 either in thick 1–100 bar pure-CH4 atmospheres, or as a trace species in a N2- or O2-dominated atmosphere, CH4 is unlikely on TRAPPIST-1 c from a modelling standpoint because of its vulnerability to destruction via photochemistry (Turbet et al., 2018).
@@
Since O2 is a heavier molecule than N2, the O2 background scenarios generally correspond to more compact, higher mean molecular weight atmospheres. Therefore, although the results are qualitatively similar to the N2-background constraints, our transmission spectrum allows for larger amounts of the tested molecular absorbers in an O2-dominated atmosphere. As such, we unfortunately are not able to rule on any of the possible scenarios presented by Zieba et al. (2023) based on their MIRI eclipse observations of this planet. The photometric eclipse depth of 42194 ppm at 15 µm was broadly consistent (better than 2) with a 0.01 bar pure CO2 atmosphere, or trace amounts of CO2 in a O2-dominated atmospheres with a surface pressure less than 10 bar. We find that such an O2-CO2 atmosphere would give rise in transmission to a 40 ppm CO2 absorption feature at 2.7 µm (at the red edge of the SOSS detector), and a larger 90 ppm absorption band at 4.3 µm. As demonstrated by Taylor et al. (2023), CO2 cannot be reliably probed at 2.7 µm with SOSS alone, even for a hot-Jupiter, whereas the 4.3 µm feature could potentially be detectable with five transits with NIRSpec/G395M or PRISM.

### v3

@@ 3.1 Transit Light Source Effect Modelling
We first model the transmission spectrum of TRAPPIST-1 c assuming that the TLSE can completely explain each visit’s transmission spectrum. This approach is motivated by the presence of small features resembling water absorption (e.g., around 1.4 µm; Figure 2) and a pronounced slope towards short wavelengths in the spectrum of the first visit. We use the open-source code stctm (Piaulet-Ghorayeb, 2024; Piaulet-Ghorayeb et al., 2024) to perform the TLS retrievals. We model the presence of two populations of heterogeneities: spots (covering a fraction of the photosphere, and cooler than the photosphere at a temperature ), and faculae (with representing their covering fraction, and their temperature).
@@
For each population of heterogeneities, we fit for the temperature difference to the photosphere (), and place a conservative uniform prior from 0 to 50% on the covering fraction for each of the components, . The stellar photosphere temperature is fitted using a Gaussian prior, with the mean set by the literature value of 2566 K (Agol et al., 2021) and a standard deviation of 50 K. We allow faculae to be at most 1000 K hotter than the photosphere (with an otherwise uniform prior on ), while the minimum temperature of spots is set by the minimum temperature for which stellar models are available in our grid (2300 K). In the retrieval, we select at each iteration the closest-matching model (in effective temperature and surface gravity space) for the photosphere, spot, and faculae components from a finely-sampled pre-computed grid obtained with the MSG module (Townsend & Lopez, 2023) using the PHOENIX (Husser et al., 2013) stellar model grid. We also introduce the fitting of (Piaulet-Ghorayeb et al., in prep), the difference between the photosphere and heterogeneity components, allowing the heterogeneity spectra to use a lower than that of the photosphere, motivated by recent work (e.g., Fournier-Tondreau et al., 2024).
@@
We use the affine-invariant Markov Chain Monte Carlo sampler emcee (Foreman-Mackey et al., 2013) to explore the parameter space, with the number of walkers set as 20 times the number of fitted parameters, and run the chains for 5000 steps, visually monitoring for convergence, and discard the first 60% as burn-in. In addition to the stellar heterogeneity properties, we fit the wavelength-independent transit depth of TRAPPIST-1 c, . We perform independent fits to the spectrum of each visit, as they are separated by about one year and should not, a priori, be affected by the same TLS contribution. We also perform additional retrievals considering only faculae or only spots in addition to the quiet stellar photosphere.
We find that when either only spots or both spots and faculae are included in our model, we obtain a very good match to the observed spectra, without the need for a planetary atmosphere. The addition of faculae enables us to fit better the shortest-wavelength data points with a tentative downward slope which cannot be created by spots alone (Figure 2) and results in broader distributions on the spot properties given the interplay with a facula component. However, when only faculae are included in our retrievals, the resulting TLS model does not capture the observed spectral features, particularly the broad absorption features around 1.4 and 1.9 µm (as faculae would predict these features to instead be inverted). The resulting parameters from these retrievals are therefore not presented, as they result in poorer fits.
We find that similar spot and facula properties can explain both visits (Figure 2; Table 2), which motivates the use of a shared stellar contamination component in our joint retrievals of stellar and planetary atmosphere contributions to the spectra in Section 4. There is, however, a slight () offset between the “bare-rock” transit depths retrieved in the fit to visit 1 ( ppm) and visit 2 ( ppm). Finally, our retrievals do not support the use of different values for the photosphere and heterogeneity components, as we only find lower limits on (e.g., Figure 8).
We use these TLS-only retrievals to produce “TLS-corrected” spectra which we use to explore the limiting effect of stellar contamination on the constraining power of our observations (see Section 5.5; Figure 9), and for visualization purposes in order to interpret the results from our atmosphere retrievals (Figures 4 and 5). For each visit, the TLS-corrected spectrum is produced by dividing the observed spectrum by the best-fitting stellar contamination model, and then multiplying the result by the best-fit model bare-rock transit depth (6776 ppm for visit 1, 6730 ppm for visit 2). These spectra aim to quantify the precision we could reach on atmospheric inferences given our observing setup if stellar contamination was not a factor.
3.2 Out-of-Transit Stellar Spectrum Modelling
In order to verify the reliability of our TLS retrievals described above, as well as gain additional insights into the heterogeneous nature of the photosphere of TRAPPIST-1, we directly model the out-of-transit stellar spectrum, roughly following the methodologies of Wakeford et al. (2019) and Moran et al. (2023). For each visit, we first create a median stack of the stellar spectrum considering only the out-of-transit integrations. For visit 1, this includes both the pre- and post-transit baseline. For visit 2, we only include integrations before the flare event in the pre-transit baseline. The post-transit baseline appears, by eye, to be at a slightly higher level than the pre-transit baseline. We test fits for visit 2 using only the pre-transit and only the post-transit baselines, as well as both and note that we do not find any difference in the inferred photosphere or heterogeneity parameters as a result of these choices, and we therefore proceed using both the pre- and pos-transit baseline for the fits.
We then flux calibrate the stellar spectra, following the procedure presented in Lim et al. (2023). Like Moran et al. (2023), we use the standard deviation of the stellar spectra as a function of time as the errors on each wavelength point, as we find this more representative than the pipeline-extracted flux errors.
As with the TLS fits, we proceed to fit one-, two-, and three-component models (i.e., photosphere-only, photosphere + spots/facule, photosphere + spots + faculae) directly to the stellar spectra using the StellarFit222https://github.com/radicamc/StellarFit package. For consistency with the in-transit fits, we explore a grid of PHOENIX model spectra (Husser et al., 2013), but also test a grid of SPHINX models (Iyer et al., 2023), which include temperatures as low as 2000 K, more appropriate for modelling cold spots on ultra-cool dwarfs like TRAPPIST-1. All models in a given grid are scaled by to represent the flux received at Earth, using =0.1192 for the stellar radius (Agol et al., 2021) and =12.1 pc for the distance of the system to the Earth (Gillon et al., 2016).
One-component fits have four free parameters: the stellar effective temperature, , the surface gravity, as well as a spectrum scaling factor (e.g., Wakeford et al., 2019; Moran et al., 2023), and a multiplicative error inflation term. The photosphere temperature was allowed to vary from 2300 – 5000 K for the PHOENIX grid or 2000 – 4000 K for the SPHINX grid, the log gravity from 3.5 – 5.5 (PHOENIX) or 4.0 – 5.5 (SPHINX), and the scale factor from 0.8 – 1.2. Multi-component fits also include the spot/faculae temperature (/), covering fraction (/), and the gravity of the heterogeneity (e.g., Lim et al., 2023; Fournier-Tondreau et al., 2024). Spot/faculae temperatures are required to be at least 100 K and up to 1000 K cooler/warmer than the photosphere, whereas their gravity must be within 1.0 of the photosphere value.
Unlike previous works, we do not simply consider a “best-fitting” model, but instead explore the full parameter space of model fits to obtain distributions on the photosphere and heterogeneity parameters analogous to our TLS retrievals. We use nested sampling for the posterior exploration, as implemented in the dynesty (Speagle, 2020) package, which also allows us to assess the statistical evidence for including heterogeneities in our models. The best fitting three-component models for both stellar grids are shown in Figure 3, and are in excellent agreement with the in-transit TLS results. The best-fitting parameters from both the in-transit and out-of-transit analyses are also presented in Table 2.
@@ 4 Joint modelling of atmosphere and stellar contamination
We perform three categories of retrievals corresponding to different assumptions about the planetary atmosphere: (1) H2/He-dominated scenarios; (2) single-component high-MMW atmospheres; (3) two-component high-MMW atmospheres with one infrared absorber present in the atmosphere and either N2 or O2 acting as the background gas.
@@
We parameterize H2/He-dominated atmospheres in the retrieval with a H2/He mixture as the background gas (assuming He/H2= 0.157), log-uniform priors on the abundances of N2, CH4, H2O, CO, CO2, and NH3, and a log-uniform prior on the pressure at which the atmosphere becomes opaque (either due to the presence of a cloud deck or a solid surface). We use the HELIOS-K (Grimm & Heng, 2015; Grimm et al., 2021) compiled line lists for CH4 (Hargreaves et al., 2020), H2O (Polyansky et al., 2018), CO (Hargreaves et al., 2019), CO2 (Yurchenko et al., 2020), and NH3 (Coles et al., 2019).
@@
For single-component high-MMW atmospheres, we only explore absorbers with prominent features over the NIRISS/SOSS wavelength range; namely, atmospheres made of 100% CH4, H2O, CO, or NH3. We also fit for the “effective surface pressure” at which the atmosphere becomes opaque.
@@
Finally, we also explore potential N2-dominated atmospheres (similar to some terrestrial objects in the solar system), or O2-dominated atmospheres (motivated by the possible O2-CO2 atmosphere consistent with the MIRI eclipse observations of TRAPPIST-1 c; Zieba et al., 2023) where only one of CH4, H2O, CO, or NH3 acts as an additional absorber beyond the broad N2-N2 or O2-O2 collision-induced absorption features. In this retrieval setup, we fit for the partial pressure of each component. Unfortunately, the CH4 line lists does not extend all the way to the short-wavelength end of NIRISS/SOSS order 2 (Hargreaves et al., 2020). Therefore, we exclude the order 2 data points from the spectrum for the 100% CH4, N2-CH4, and O2-CH4 retrievals.
@@
We assume that the vertical temperature structure is isothermal, motivated by the fact that our transmission spectra would not the precision to measure any temperature gradient in the thin transit photosphere region we are probing. We use the nested sampling method (Skilling, 2004, 2006) to sample the full parameter space, with the multi-ellipsoid method implemented using the nestle module333https://github.com/kbarbary/nestle within the SCARLET retrieval framework. The final posterior distributions are obtained from retrievals performed using at least 1000 live points. We compute forward models at R=31,250 before convolving them to the resolution of the spectrum, assuming a uniform throughput within each bin for the likelihood calculation.
@@
We also calculate forward models for 1 and 100 solar metallicity, H2/He-dominated atmospheres with varying cloud-top pressures (Figure 4; left panel), and models for several high MMW compositions (Figure 5; top panel), to put our retrieval results in context. For each model, the temperature-pressure profile is assumed to be isothermal at the equilibrium temperature of TRAPPIST-1 c.
For each set of retrievals we account for the impacts of stellar contamination, thereby marginalizing over the interplay between unocculted stellar heterogeneities and planetary absorption features. In addition to the planetary atmosphere properties, Motivated by the results in Section 3, we assume two heterogeneity populations: spots (cooler than the photosphere), and faculae (hotter than the photosphere). For each heterogeneity component, we fit a covering fraction (with a wide uniform prior of 0 to 50%) and a temperature contrast to the photosphere (following the uniform priors of the TLS-only fit). We also fit for the photosphere temperature itself, with a Gaussian prior based on the value reported by Agol et al. (2021).
We allow the photosphere to vary within the range as lower values are required to obtain a good match of the PHOENIX models to the out-of-transit spectrum of TRAPPIST-1 (see e.g., Lim et al., 2023 and Section 3.2). Given that the TLS-only fit did not provide evidence for a difference in between the photosphere and heterogeneity components, we fix them to the same value for each model evaluation. Finally, we fit for an offset between the visit 1 and visit 2 spectra, motivated by the stctm retrieval results.
Given the consistency between the stellar heterogeneity parameters inferred from both the TLS and stellar spectrum fits, we assume that the heterogeneity components for both visits are identical in our joint atmosphere fits. However, we note that we also performed one set of atmosphere+stellar contamination retrievals where we fit for visit-specific stellar heterogeneity properties, and found that this does not affect our inferences. We show the results for H2/He-dominated atmospheres in Figure 4, and high-MMW atmospheres in Figures 5 and 6.
@@ 5 Discussion
5.1 Inferences of the Stellar Photosphere
For the out-of-transit stellar modelling using both grids and visits, the photosphere temperature is generally higher than, but still broadly consistent with literature values (e.g., Agol et al., 2021; Davoudi et al., 2024). The retrieved log gravities (3.8 for PHOENIX and 4.2 for SPHINX) are both significantly lower than the 5.2 found by Agol et al. (2021) and Davoudi et al. (2024), through this discrepancy was also noted in the TRAPPIST-1 spectra obtained with NIRISS/SOSS by Lim et al. (2023) in their observations of planet b.
Our results from both grids are broadly consistent with the TLS retrievals (e.g., Figures 2, 3; Table 2). For SPHINX, a two-component model including a photosphere and spots is preferred at 4 (using the Benneke & Seager (2013) Bayesian evidence to -level conversion) over a homogeneous photosphere model and at 2 over a model including both spots and faculae for both visits. Like in the TLS retrievals, when faculae are included, the fits push the covering fraction to zero in order to minimize their contribution. Both the spot covering fraction and temperature are in excellent agreement with the TLS results. The spot temperatures are also consistent with those inferred by previous studies (e.g., Wakeford et al., 2019; Garcia et al., 2022; Lim et al., 2023; Davoudi et al., 2024). For the PHOENIX grid, the results are also in excellent agreement with both the SPHINX grid and the TLS fit results (Table 2). In this case, the three-component fit is preferred (though the significance is low; 1.8). The faculae covering fraction is also only 3%, and therefore essentially negligible. For all grids tested, we find that the gravity of the heterogeneities is always consistent with that of the photosphere.
We note though that the overall quality of the fits are unsatisfactory, with values generally 350. The short wavelengths (1.1 µm) are especially badly fit by the PHOENIX grid in particular. In fact, the PHOENIX grid struggled to fit both the short and long wavelengths simultaneously, routinely overestimating the flux 1.7 µm and finding photosphere temperatures that were highly inconsistent with literature values. We therefore restricted the models to wavelengths 1.1 µm for the PHOENIX grid where the models were better matches to the data. Performing the same test for the SPHINX grid does not alter the retrieved parameter distributions.
Another caveat is that the high resolution of NIRSpec/G395H (2700) allowed Moran et al. (2023) to fit to individual stellar lines. However, in our case, the medium resolution of NIRISS/SOSS (700) does not provide this same capability (see the inset in Figure 3), and our fits are more sensitive to the broadband structures in the spectrum. In all, though we confirm the results of the in-transit TLS fits, our stellar spectrum analysis demonstrates the clear need for improved stellar models for ultra-cool stars. Moreover, we posit that since the in-transit analysis relies on ratios of stellar models instead of absolute fits, it may be more robust to the aforementioned model inaccuracies than direct fits to stellar spectra — something that should be explored further.
@@ 5.2 Cloud-Free and Thick Hydrogen-Rich Atmospheres Are Ruled Out
We find that, similarly to the closer-in planet TRAPPIST-1 b (Lim et al., 2023), we can confidently rule out thick or cloud-free hydrogen-dominated atmospheres, with a 3- limit of 10 mbar on the effective surface pressure of a H2/He-dominated atmosphere atop TRAPPIST-1 c (Figure 4). The thin hydrogen-rich atmosphere scenarios allowed by the transmission spectrum are, moreover, unlikely from an atmospheric evolution standpoint — TRAPPIST-1 c was exposed to large cumulative amounts of high-energy XUV irradiation from its host star, sufficient to strip away any primordial atmosphere (Hori & Ogihara, 2020; Turbet et al., 2020) even when considering hydrogen outgassing (Hu et al., 2023). The transmission spectrum still allows within the 3- confidence contours for thick, high-metallicity (100 solar) atmospheres, which may be more resilient against escape due to their higher MMWs.
5.3 Limits on High-Mean Molecular Weight Atmospheres
@@
We, therefore, turn to evaluate the range of plausible high-metallicity atmosphere scenarios for TRAPPIST-1 c, and focus on the detectability of H2O-, NH3-, CO- and CH4 in high-MMW atmospheres, as these infrared absorbers have sufficient opacity within the NIRISS/SOSS wavelength range to enable meaningful constraints.
As expected from the high-metallicity tail of the surface pressure-metallicity distribution (Figure 4), the densest atmosphere cases we tested do not provide strong upper limits on the effective surface pressure. We find that deep, 100% H2O, NH3, and CH4 atmospheres remain compatible with our spectrum at the 3 level, although 1 bar atmospheres are disfavored at the 2 level (Figure 5). In the 100% CO case, we exclude even a 1-bar atmosphere at the 3 level. However, we caution that this result mostly stems from the fact that CO has minimal opacity in the SOSS waveband from 2–2.5 µm, and a pure-CO atmosphere would elsewhere yield a perfectly flat spectrum. Therefore, this example only rules out 100% CO atmospheres, but likely not cases where a CO-rich atmosphere is mixed with small amounts of another infrared absorber.
@@
We then turn to models that allow for lower atmospheric MMWs, where the infrared absorber can be a trace species rather than a dominant component — representative of the atmospheres of some solar system planets. In this scenario, the background gas is N2 or O2, with only one of the aforementioned species as the active infrared absorber (Figure 6). We compare these limits, where available, with measured compositions of solar system bodies. In particular, we highlight in Figure 6 the amount of H2O in Earth’s N2-dominated atmosphere (using 0.4% as the mean mole fraction over the entire 1-bar atmosphere, although it can reach about one percent at sea level; Kasai et al., 2011; Kelsey et al., 2022), or the amount of CH4 in Titan’s atmosphere, assuming either a cloud-free Titan scenario (where one can see down to the surface) with 5% CH4, or a hazy scenario accounting for the fact that hydrocarbon haze opacity raises the effective surface pressure to 10 mbar, and with a factor of three less methane in the upper atmosphere (see e.g. Charnay et al., 2014).
We find that similar amounts of water to what is observed in Earth’s atmosphere are disfavored at the 2 level for TRAPPIST-1 c, while the transmission spectrum alone is not sensitive enough to the CH4 abundance to offer meaningful constraints on Titan-like atmospheres. Although our observations cannot rule out atmospheres where NH3 or CH4 make up the entire volatile budget (or are trace species in N2- or O2-dominated atmospheres), the presence of methane and ammonia is unlikely on TRAPPIST-1 c from a modelling standpoint because of its vulnerability to destruction via photochemistry (Turbet et al., 2018).
@@
Since O2 is a heavier molecule than N2, the O2 background scenarios generally correspond to more compact, higher-MMW atmospheres. Therefore, although the results are qualitatively similar to the N2-background constraints, our transmission spectrum allows for larger amounts of the tested molecular absorbers in an O2-dominated atmosphere. Further, because of the lack of sensitivity to CO2 absorption with NIRISS/SOSS spectroscopy, we are not able to rule on any of the possible scenarios presented by Zieba et al. (2023) based on their MIRI eclipse observations of this planet. The photometric eclipse depth of 42194 ppm at 15 µm was broadly consistent (better than 2) with a 0.01 bar pure CO2 atmosphere, or trace amounts of CO2 in a O2-dominated atmospheres with a surface pressure less than 10 bar. We find that such an O2-CO2 atmosphere would give rise in transmission to a 40 ppm CO2 absorption feature at 2.7 µm (at the red edge of the SOSS detector), and a larger 90 ppm absorption band at 4.3 µm. As demonstrated by Taylor et al. (2023), CO2 cannot be reliably probed at 2.7 µm with SOSS alone, even for a hot-Jupiter, whereas the 4.3 µm feature could potentially be detectable with five transits with NIRSpec/G395M or PRISM.